# thanks: These two authors contributed equally

### v1

@@ IV.3 External illumination: anti-Bragg
Figure 8: Spectrum of the incoherent component $\left|\psi_{\mathrm{incoh}}(\omega)\Gamma/\alpha_{\mathrm{sc}}^{2}\right|$ for atoms illuminated under the anti-Bragg angle condition and $\Delta=0$. The curves are obtained for $\beta=0.01$ and different atom numbers. For small $N$, a coherent built-up takes place, while for large $N$ the spectrum saturates to $\alpha^{2}_{\mathrm{sc}}/(\beta(1-\beta)$ as shown by the black dashed line. For comparison, the dotted gray line shows $\left|\psi_{\mathrm{incoh}}(\omega)\Gamma/\alpha_{\mathrm{sc}}^{2}\right|$ for illumination through a waveguide [2] (or equivalently in Bragg configuration) for the same $\beta$ and the atom number yielding to the maximal value at $\omega=0$.
Figure 9: a) Incoherent component of the two-photon wavefunction $|\psi_{\mathrm{incoh}}(\tau=0)|$ as a function of emitter $N$. In contrast to the Bragg/transmission illumination $|\psi_{\mathrm{incoh}}(\tau=0)|$ grows much slower. For comparison, the two-photon field is normalized to the scattered coherent field $\alpha_{\mathrm{sc}}^{2}$ of the first emitter in both cases. b) The second-order correlation function $g^{(2)}(0)$ as a function of emitters $N$ in solid blue for the anti-Bragg configuration. For comparison the transmission configuration is shown in orange and the Bragg configuration in green, where $g^{(2)}(0)$ quickly reaches $1$ after an initial antibunching for $N=1$ (shown in the inset zoom). Both figures are shown for $\beta=0.01$.
Now we look at the anti-Bragg configuration [76], i.e., the case where emitted coherent fields interfere destructively inside the waveguide, in contrast to the Bragg case. At resonance, the anti-Bragg angle $\zeta$ is given by:
@@
which can be simplified by a geometric sum to
@@
In contrast to the above case studies, the driving Rabi frequency remains constant along the chain, which leads to different spectral properties of the incoherent component. In Fig. 8 we show $|\psi_{\mathrm{incoh}}|$. It is apparent that, for large enough $N$, such that $|t_{\omega=0}|^{N}\simeq 0$, $\psi_{\mathrm{incoh}}(\omega)$ exhibits a plateau around $\omega=0$, which asymptotically reaches a value of $-\alpha_{\mathrm{sc}}^{2}\Gamma/(\beta(1-\beta))$, as indicated by the black dashed line in Fig. 8.
The strong incoherent amplitude $\psi_{\mathrm{incoh}}(\omega)$ can be directly linked to squeezed light via Eq. (6), and we obtain a maximal squeezing of
@@
within the plateau. We note that Eq. (38) is only valid for $\Omega^{\mathrm{ext}}\ll\Gamma$. Specifically, it should not be concluded that perfect squeezing (i.e., $S_{0}(\omega)=-1/4$) can be obtained.
To quantify the strength of the non-linearity, we integrate $\psi_{\mathrm{incoh}}(\omega)$ over all frequencies, $\psi_{\mathrm{incoh}}(\tau=0)=\int_{-\infty}^{\infty}\frac{\mathrm{d}\omega}{2\pi}\psi_{\mathrm{incoh}}(\omega)$. This quantity can be experimentally inferred from the total measured noise reduction [2].
To compare different configurations, we normalize $\psi_{\mathrm{incoh}}$ to the scattering rate of the first emitter $\left(\alpha_{1}^{\mathrm{sc}}\right)^{2}$. With this normalization, $\psi_{\mathrm{incoh}}$ becomes independent of the input power and the three different configurations are compared in Fig. 9. From Fig. 9 a) it is apparent that $|\psi_{\mathrm{incoh}}(\tau=0)|$ reaches much larger values in the anti-Bragg configuration than under illumination through the waveguide (or, equivalently, for the Bragg-configuration). This larger amount of incoherent photon pairs is due to the fact that the driving strength along the chain does not decrease, in contrast to the exponentially damped driving in both Bragg illumination and illumination through the waveguide.
Let us finally discuss the normalized second-order correlation function $g^{(2)}(\tau)$ in the anti-Bragg configuration, see Fig. 9 b). In our approximation of Eq (4), it diverges for every even number of atoms, since the coherent power $P_{\mathrm{coh}}\propto|\alpha_{\mathrm{out}}|^{2}$ is zero. For odd atom numbers, one obtains:
@@
While the coherent two-photon component is independent of $N$ with a value of $\alpha^{2}_{\mathrm{sc}}$, the incoherent two-photon component $\psi_{\mathrm{incoh}}(\tau=0)$ grows much faster than in the two other configurations. This leads to strong bunching already for much smaller atom numbers that when illuminating through the waveguide, see Fig. 9 b). Thus, the anti-Bragg configuration is promising for solid state emitters, where obtaining large ensembles of identical emitters is challenging. For example, already for two emitters, perfect bunching (i.e., diverging $g^{(2)}(\tau=0)$) is predicted [78]. Note that, the underlying destructive interference could be even obtained for randomly spaced emitters with inhomogenous transition strengths and frequencies. This only requires individual addressability including the ability to accordingly adapt the local phases and amplitudes of the external drive, e.g., with a spatial light modulator (SLM). The increased yield of correlated photon pairs could then be of practical use for heralded single photon sources.
@@ IV.4 Combined illumination through waveguide and external
In this last configuration, we study the case where many atoms are driven both through the waveguide and externally as sketched in Fig. 10. As discussed in Sec. III.3 for a single atom and $\beta\ll 1$, the system effectively behaves as if it was solely illuminated through the waveguide with a rescaled coupling constant $\beta^{\prime}=\beta\left(1+\Omega^{\mathrm{ext}}/\Omega^{\mathrm{wg}}\right)$.
For many emitters, we restrict our discussion to a resonant excitation ($\Delta=0$) with both fields in phase. Then, this concept of an effectively increased coupling constant greatly simplifies the problem. Specifically, the output coherent photon amplitude is given by
@@
where the individual amplitude transmission coefficients are given by $t^{\prime}(n)=1-2\beta^{\prime}(n)$. Here, the coupling strength of the n-th emitter is given by $\beta^{\prime}(n)/\beta=1+\Omega^{\mathrm{ext}}/(\Omega_{n}^{\mathrm{wg}}t^{n-1}_{0})$.
For the incoherent component, the creation process behaves as if each coupling constant $\beta^{\prime}(n)$ was enhanced. However, the absorption of the incoherent photon pairs $|t_{\omega}|^{2(N-n)}$ continues to scale with the bare value of $\beta$ This leads to:
@@
where $|\alpha_{\mathrm{in}}|^{2}$ is the power sent through the waveguide, and $g_{\omega}^{\prime}(n)=\frac{2\beta^{\prime}(n)}{1-2i\omega/\Gamma}$ the position dependent photon generation coefficient. In analogy to $t^{\prime}(n)$, we obmit the subscript for $g^{\prime}(n)$, as here we assumed $\Delta=0$.
@@
Figure 11: The second-order correlation function $g^{(2)}(0)$ as function of the numbers of emitters in combined illumination under the Bragg angle and through a waveguide for $\Delta=0$ for three exemplary parameter sets. For comparison the black dashed line shows the case of pure waveguide illumination with a bare coupling constant $\beta=0.01$. By adding the external Bragg beam with a field strength $\Omega^{\mathrm{ext}}/\Omega^{\mathrm{wg}}_{\mathrm{in}}=2$ at the first emitter, the collective non-linear interaction of the ensemble is enhanced such that for $N=29$ a point of perfect antibunching is reached (orange line). For $\Omega^{\mathrm{ext}}/\Omega^{\mathrm{wg}}_{\mathrm{in}}=10$ this point is already reached with 6 atoms. When increasing the coupling strength this effect becomes more pronounced, and with $\beta=0.03$ antibunching can be reached with three emitters and $\Omega^{\mathrm{ext}}/\Omega^{\mathrm{wg}}_{\mathrm{in}}=5$. The solid lines are guides to the eyes.
@@ V Conclusion & Summary
In summary, we have presented a theoretical model to efficiently calculate the nonlinear response of an ensemble of $N$ quantum emitters weakly coupled ($\beta\ll 1$) to a single propagating optical mode in the low saturation regime ($\Omega\ll\Gamma$). Using this model, we derive simple analytic expressions for the incoherent component of the two-photon wavefunction, $\psi_{\mathrm{incoh}}(\omega)$, for numbers of emitters for which alternative state-of-the-art methods [33, 60] are numerically costly. This allows us to infer both the squeezing spectrum and the second-order correlation function. We show how large ensembles, as well as countable numbers of identical weakly-coupled emitters can produce quantum states of light with sub/super Poissonian statistics and squeezing. While this is technologically achievable with cold atoms coupled to nanofibers as demonstrated in [1, 2, 3], obtaining large ensemble of identical emitters is challenging with solid state systems. To circumvent this limitations, we present two techniques that allow one to arbitrarily reduce the number of emitters required to observe these strong quantum effects.
It is an open question how the discussed effects extends to larger saturation. Recently, on-resonant illumination through the waveguide was studied also for larger saturation parameters [60] but remains unexplored in the off-resonant case, which is also covered by our model. Furthermore, our model could be extended such that it also yields reliable predictions for larger values of $\beta$. In this regime, interesting effects such as photon number sorting [79], topologically protected edge states [80, 81], and many-body-localized phases [82] have been predicted. Here, we expect our model to yield physical insight in the case of large ensembles where other approaches exclusively offer numerical predictions.
@@ Appendix A Unidirectional coupling
In this article, we assume a perfectly chiral system where all emitters scatter only in the forward direction within the waveguide. This assumption is motivated by several factors: First, such chirality exits naturally in setups based on cold atoms coupled to nanofibers [69]. Second, unidirectional coupling drastically reduces the complexity of the calculation, and third, for illumination through the waveguide we expect only small differences between chiral and symmetric coupling [33]. The same is expected to hold for the anti-Bragg configuration. However, in the Bragg configuration the backward component interferes constructively, leading to a strong backward component that can no longer be neglected, and we expect that our model to work best with directional coupling.
@@
Figure 12: a) The incoherent component $\psi_{\mathrm{incoh}}(\tau=0)$ for $\beta=0.01$ and different detuning $\Delta$. In the off-resonant case $\psi_{\mathrm{incoh}}(\tau=0)$ shows a complicated dynamic. For small number of emitters (low OD-regime), it first increases linearly with $N$ due to the coherent enhancement. For larger $N$ a more complicated picture emerges due to interferences between the light scattered by different emitters that can even lead to a cancellation of $\psi_{\mathrm{incoh}}(\tau=0)$. The dotted lines show the low OD approximation of coherent build-up in forward scattering. b) Evolution of the incoherent and coherent component at zero delay, as well as the corresponding $g^{(2)}(\tau=0)$ as a function of $N$ for the same $\beta$ and $\Delta=0$. At $N=150$ the point of destructive interference occurs and for larger $N$ strong bunching can be observed [33, 1]. The dotted line is the further approximation of Eq. 45 with good agreement for small $\beta$.
@@
For a larger number of emitters, Eq. (43) predicts sidebands which develop due to the absorption of the incoherent photon pairs from the ensemble of emitters and which have been observed in [2]. Eq. (43) has a divergence for $\omega=0$. This divergence can be removed by using $1+4\omega^{2}/\Gamma^{2}\simeq\omega^{2}\Gamma^{2}$. This is a good approximation in the large OD regime where most of the incoherent photons are in the sidebands and around $\omega\simeq 1$ the weight of $\psi_{\mathrm{incoh}}(\omega)$ is small. From this we can calculate:
@@
Note that, Eq. (45) represents a very good approximation for $g^{(2)}(0)$ even for a small OD, as shown in Fig. 12. However, this is a coincidence due to the fact that in the low $OD$-limit, $g^{(2)}(0)$ is exponentially dominated by the coherent component $\alpha_{\mathrm{in}}t_{\Delta}^{2N}$ and the error on $\psi_{\mathrm{incoh}}(\tau=0)$ has a negligible influence. A similar expression was found in [13] based on a Green’s function approach.

### v2

@@ IV.3 External illumination: anti-Bragg
Figure 8: Spectrum of the incoherent component $\left|\psi_{\mathrm{incoh}}(\omega)\Gamma/\alpha_{\mathrm{sc}}^{2}\right|$ for atoms illuminated under the anti-Bragg angle condition and $\Delta=0$. The curves are obtained for $\beta=0.01$ and different atom numbers. For small $N$, a coherent built-up takes place, while for large $N$ the spectrum saturates to $\alpha^{2}_{\mathrm{sc}}/(\beta(1-\beta)$ as shown by the black dashed line. For comparison, the dotted gray line shows $\left|\psi_{\mathrm{incoh}}(\omega)\Gamma/\alpha_{\mathrm{sc}}^{2}\right|$ for illumination through a waveguide [2] (or equivalently in Bragg configuration) for the same $\beta$ and the atom number that maximizes $\left|\psi_{\mathrm{incoh}}(\omega)\right|$ at $\omega=0$.
Figure 9: a) Incoherent component of the two-photon wavefunction $|\psi_{\mathrm{incoh}}(\tau=0)|$ as a function of the number of emitters $N$. In contrast to the Bragg/transmission illumination $|\psi_{\mathrm{incoh}}(\tau=0)|$ grows much faster in the anti-Bragg configuration. For comparison, the two-photon field is normalized to the scattered coherent field $\alpha_{\mathrm{sc}}^{2}$ of the first emitter in both cases. b) The second-order correlation function $g^{(2)}(0)$ as a function of emitters $N$ in solid blue for the anti-Bragg configuration and for odd $N$. For comparison, $g^{(2)}(\tau=0)$ for the transmission configuration is shown in orange and for the Bragg configuration in green, where $g^{(2)}(0)$ quickly reaches $1$ after an initial antibunching for $N=1$ (shown in the inset zoom). Both figures are calculated for $\beta=0.01$.
Now we look at the anti-Bragg configuration [78], i.e., the case where emitted coherent fields interfere destructively inside the waveguide, in contrast to the Bragg case. Again, we restrict our discussion to the resonant case and the anti-Bragg angle $\zeta$ is given by:
@@
this geometric sum can be simplified to
@@
In contrast to the above case studies, the driving Rabi frequency remains constant along the chain, which leads to a continuous build-up of the incoherent component with $N$ and different spectral properties of the incoherent component. In Fig. 8 we show $|\psi_{\mathrm{incoh}}(\omega)|$. It is apparent that, for large enough $N$, such that $|t_{\omega=0}|^{N}\simeq 0$, $\psi_{\mathrm{incoh}}(\omega)$ exhibits a plateau around $\omega=0$. This plateau asymptotically reaches a value of $-\alpha_{\mathrm{sc}}^{2}\Gamma/(\beta(1-\beta))$, as indicated by the black dashed line in Fig. 8.
The strong incoherent component $\psi_{\mathrm{incoh}}(\omega)$ can be directly linked to squeezed light via Eq. (7), and we obtain a maximal squeezing of
@@
within the plateau and we note that Eq. (39) is only valid for $\Omega^{\mathrm{ext}}\ll\Gamma$  [^10] Specifically, it should not be concluded that perfect squeezing (i.e., $S_{0}(\omega)=-1/4$) can be achieved as our model loses validity when $\Omega_{\mathrm{ext}}\simeq\Gamma$..
To quantify the strength of the atom’s non-linearity, we integrate $\psi_{\mathrm{incoh}}(\omega)$ over all frequencies, $\psi_{\mathrm{incoh}}(\tau=0)=\int_{-\infty}^{\infty}\frac{\mathrm{d}\omega}{2\pi}\psi_{\mathrm{incoh}}(\omega)$. This quantity can be experimentally inferred from the noise reduction in the amplitude quadrature [2].
To compare different configurations, we normalize $\psi_{\mathrm{incoh}}(\tau)$ to the scattering rate of the first emitter $\left(\alpha_{1}^{\mathrm{sc}}\right)^{2}$. With this normalization, $\psi_{\mathrm{incoh}}(\tau)$ becomes independent of the input power and the three different configurations are compared in Fig. 9. From Fig. 9 a) it is apparent that for the same number of atoms $|\psi_{\mathrm{incoh}}(\tau=0)|$ reaches much larger values in the anti-Bragg configuration than under illumination through the waveguide (or, equivalently, for the Bragg-configuration). This larger amount of incoherent photon pairs is due to the fact that the driving strength along the chain does not decrease, in contrast to the exponentially damped driving in both Bragg illumination and illumination through the waveguide.
Let us finally discuss the normalized second-order correlation function $g^{(2)}(\tau)$ in the anti-Bragg configuration, see Fig. 9 b). In our approximation of Eq (5), it diverges for every even number of atoms, since the coherent power $P_{\mathrm{coh}}\propto|\alpha_{\mathrm{out}}|^{2}$ is zero  [^11] Note that this is not a limitation of our model, as this divergence in $g^{(2)}(0)$ strictly only occurs for vanishing input power. For finite input power high-order corrections will lead to finite photon bunching.. For odd atom numbers, one obtains:
@@
While the coherent two-photon component is independent of $N$ with a value of $\alpha^{2}_{\mathrm{sc}}$, the incoherent two-photon component $\psi_{\mathrm{incoh}}(\tau=0)$ grows much faster than in the two other configurations. This leads to strong bunching already for much smaller atom numbers that when illuminating through the waveguide, see Fig. 9 b). Thus, the anti-Bragg configuration is promising for solid state emitters, where obtaining large ensembles of identical emitters is challenging. For example, already for two emitters, perfect bunching (i.e., diverging $g^{(2)}(\tau=0)$) is predicted [82]. Note that, the underlying destructive interference could be even obtained for randomly spaced emitters with inhomogeneous transition strengths and frequencies. This only requires individual addressability including the ability to accordingly adapt the local phases and amplitudes of the external drive, e.g., with a spatial light modulator (SLM). The increased yield of correlated photon pairs could then be of practical use for narrowband photon pair sources, or heralded single photon sources.
@@ IV.4 Combined illumination through waveguide and external
In this last section, we study the case where many atoms are driven simultaneously via the waveguide and an external field as sketched in Fig. 10 with both fields in phase at the atoms. As discussed in Sec. III.3 for a single atom and $\beta\ll 1$, the light in the fiber effectively behave as the atoms were solely illuminated through the waveguide with a rescaled coupling constant $\beta^{\prime}=\beta\left(1+\Omega^{\mathrm{ext}}/\Omega^{\mathrm{wg}}\right)$.
Then, this concept of an effectively increased coupling constant greatly simplifies the problem. Specifically, the output coherent photon component for $N$ emitters and resonant drive ($\Delta=0$) is given by
@@
where the individual amplitude transmission coefficients are given by $t^{\prime}(n)=1-2\beta^{\prime}(n)$. Here, the coupling strength of the n-th emitter is given by
$\beta^{\prime}(n)/\beta=1+\Omega^{\mathrm{ext}}/(\Omega_{n}^{\mathrm{wg}}t^{n-1}_{0}).$ (42)
For the incoherent component, the creation process behaves as if each coupling constant was $\beta^{\prime}(n)$. However, the transimission of the incoherent photon pairs $|t_{\omega}|^{2(N-n)}$ continues to scale with the bare value of $\beta$ This leads to:
@@
where $|\alpha_{\mathrm{in}}|^{2}$ is the power sent through the waveguide, and $g_{\omega}^{\prime}(n)=\frac{2\beta^{\prime}(n)}{1-2i\omega/\Gamma}$ the position dependent photon generation coefficient. In analogy to $t^{\prime}(n)$, we omit the subscript for $g^{\prime}(n)$, as here we assumed $\Delta=0$.
@@
Figure 11: The second-order correlation function $g^{(2)}(0)$ as function of the numbers of emitters in combined illumination under the Bragg angle and through the waveguide for $\Delta=0$ for three exemplary parameter sets. For comparison, the black dashed line shows the case of pure waveguide illumination with a bare coupling constant $\beta=0.01$. By adding the external Bragg beam with a field strength $\Omega^{\mathrm{ext}}/\Omega^{\mathrm{wg}}_{\mathrm{in}}=2$ at the first emitter, the collective non-linear interaction of the ensemble is enhanced such that for $N=29$ a point of perfect antibunching is reached (orange line). For $\Omega^{\mathrm{ext}}/\Omega^{\mathrm{wg}}_{\mathrm{in}}=10$ this point is already reached with 6 atoms. When increasing the coupling strength this effect becomes more pronounced, and with $\beta=0.03$ perfect antibunching can be reached with three emitters and $\Omega^{\mathrm{ext}}/\Omega^{\mathrm{wg}}_{\mathrm{in}}=5$. The solid lines are guides to the eyes.
@@ V Conclusion & Summary
In summary, we have presented a theoretical model to efficiently calculate the nonlinear response of an ensemble of $N$ quantum emitters weakly coupled ($\beta\ll 1$) to a single propagating optical mode in the low saturation regime ($\Omega\ll\Gamma$). Using this model, we derive simple analytic expressions for the incoherent component of the two-photon wavefunction, $\psi_{\mathrm{incoh}}(\omega)$, for numbers of emitters for which alternative state-of-the-art methods [34, 62] are numerically costly. This allows us to infer both the squeezing spectrum and the second-order correlation function. We show how large ensembles, as well as countable numbers of identical weakly-coupled emitters can produce quantum states of light with sub/super Poissonian statistics and squeezing. While this is technologically achievable with cold atoms coupled to nanofibers as demonstrated in [1, 2, 3], obtaining large ensemble of identical emitters is challenging with solid state systems. To circumvent this limitation, we present two techniques that allow one to arbitrarily reduce the number of emitters required to observe these strong quantum effects.
It is an open question how the discussed effects extend to larger saturation. Recently, on-resonant illumination through the waveguide was studied also for larger saturation parameters [62] but remains unexplored in the off-resonant case, which is also covered by our model. Furthermore, our model could be extended such that it also yields reliable predictions for larger values of $\beta$. In this regime, interesting effects such as photon number sorting [83], topologically protected edge states [84, 85], and many-body-localized phases [86] have been predicted. Here, we expect our model to yield physical insights in the case of large ensembles where other approaches exclusively offer numerical predictions.
@@ Appendix A Unidirectional coupling
In this article, we assume a perfectly chiral system where all emitters scatter only in forward direction of the waveguide. This assumption is motivated by several factors: First, such chiral scattering exits naturally in setups based on cold atoms coupled to nanofibers [87]. Second, unidirectional coupling drastically reduces the complexity of the calculation, and third, for illumination through the waveguide we expect only small differences between chiral and symmetric coupling [34]. The same is expected to hold for the Bragg configuration and anti-Bragg configuration.
@@
Figure 12: a) The incoherent component $\psi_{\mathrm{incoh}}(\tau=0)$ for $\beta=0.01$ and different detuning $\Delta$ as a function of the emitter $N$. In the off-resonant case $\psi_{\mathrm{incoh}}(\tau=0)$ shows a complicated dynamic. For small number of emitters (low OD-regime), it first increases linearly with $N$ due to the coherent enhancement. For larger $N$ a more complicated picture emerges due to interferences between the light scattered by different emitters that can even lead to a cancellation of $\psi_{\mathrm{incoh}}(\tau=0)$. The dotted lines show the low OD approximation of coherent build-up in forward scattering. b) Evolution of the incoherent and coherent component at zero delay, as well as the corresponding $g^{(2)}(\tau=0)$ as a function of $N$ for the same $\beta$ and $\Delta=0$. At $N=150$ the point of destructive interference occurs and for larger $N$ strong bunching can be observed [34, 1]. The dotted line is the further approximation of Eq. 47 with good agreement for small $\beta$.
@@
For a larger number of emitters, Eq. (45) predicts sidebands which develop due to the absorption of the incoherent photon pairs from the ensemble of emitters and which have been observed in [2]. Note that Eq. (45) has a divergence for $\omega=0$. To simplify Eq. (45) one can introduce the approximation $1+4\omega^{2}/\Gamma^{2}\simeq\omega^{2}\Gamma^{2}$ in the exponent. This is a good approximation in the large OD regime where most of the incoherent photons are in the sidebands and around $\omega/\Gamma\simeq 1$ the weight of $\psi_{\mathrm{incoh}}(\omega)$ is small. From this we can calculate:
@@
Note that, Eq. (47) represents a very good approximation for $g^{(2)}(0)$ even for a small OD, as shown in Fig. 12. This is due to the fact that in the low $OD$-limit, $g^{(2)}(0)$ is exponentially dominated by the coherent component $\alpha_{\mathrm{in}}t_{\Delta}^{2N}$ and the error on $\psi_{\mathrm{incoh}}(\tau=0)$ has a negligible influence. A similar expression was found in [13] based on a Green’s function approach.